# Convergence Analysis of Non-Strongly-Monotone Stochastic Quasi-Variational Inequalities

### v2

@@ 6 Numerical Experiment
In this experiment, we define $\ell^{\text{val}}_{i}(x_{i})\triangleq\frac{1}{2}\|A_{i}^{\text{val}}x_{i}-b_{i}^{\text{val}}\|^{2},$ where $A_{i}^{\text{val}}\in\mathbb{R}^{n\times d}$ and $b_{i}^{\text{val}}\in\mathbb{R}^{n\times 1}$, and $\ell^{\text{tr}}(x)\triangleq\frac{1}{2}\|A^{\text{tr}}\mathbf{x}-b^{\text{tr}}\|^{2},$ where $A^{\text{tr}}\in\mathbb{R}^{Nn\times Nd}$ and $b^{\text{tr}}\in\mathbb{R}^{Nn\times 1}$ and $X_{i}=\{x_{i}\mid\|x_{i}\|\leq\lambda\}$ for some $\lambda>0$. One can show that this problem can ve formulated as (SQVI) by choosing $K(x)=\Pi_{i=1}^{N}K_{i}(x_{i},x_{(-i)})$, where
$K_{i}(x_{i},x_{(-i)})=\mbox{argmin}_{x_{i}\in X_{i}}\frac{1}{2}\left\|A_{i}^{\text{tr}}x_{i}+A_{(-i)}^{\text{tr}}x_{(-i)}-b_{i}^{\text{tr}}\right\|^{2},$
@@
Note that, operator $F$ is monotone and satisfies quadratic growth property but it may not be strongly monotone. In Figure 1, we present a performance comparison of our proposed methods. For the triazines dataset [7], we set the number of players $N=6$, for the eunite2001 dataset [7] $N=4$, and for the synthetic dataset $N=10$. In all cases, we utilized $80\%$ of the data points for training and allocated the remaining $20\%$ for validation. To solve the projection inexactly, observe that the sub-problem is a simple bilevel optimization problem. This type of problem has been explored in the literature [17, 38]. Here, following [38], we employed the FISTA algorithm [4] to solve the corresponding regularized problem satisfying Assumption 4. For all the experiments, we execute the inner algorithm for $k\log^{2}(k+1)(1-1\text{e-}3)^{k}$ iterations, and the remaining parameters are selected according to the following table after fine-tuning.
@@ Appendix A Extra Gradient Method
Proof of Corollary 2. (i) By taking expectation from both sides of 5, choosing $N_{k}=N$, and using Assumption 2, the following holds.
@@ Appendix B Gradient Method
Proof of Theorem 2. For any $k\geq 0$, we define $\bar{x}_{k}\triangleq\mathcal{P}_{X^{*}}(x_{k})$ where $X^{*}$ denotes the set of optimal solutions of problem (SQVI). From Lemma 2 we conclude that $\bar{x}_{k}=\mathcal{P}_{K(\bar{x}_{k})}[\bar{x}_{k}-\eta F(\bar{x}_{k})]$. Using the update rule of $x_{k+1}$, in Algorithm 2 and the fact that $e_{k}$ denotes the error of computing the projection operator, we obtain the following.
@@
Proof of Corollary 4. Taking expectation from both sides of 10, choosing $N_{k}=\lceil\rho^{-2k}\rceil$, and using Assumption 2, one can obtain:
@@
According to the assumption 4, Algorithm $\mathcal{M}$ has a convergence rate of $C/t_{k}^{2}$ within $t_{k}$ inner steps. By selecting $t_{k}=\tfrac{({k+1})\log^{2}(k+2)}{\rho^{k}}$, we conclude that $\mathbb{E}[\|{e_{k}}\|]\leq\tfrac{C}{t_{k}}=\tfrac{C\rho^{k}}{(k+1)\log^{2}(k+2)}$. By using the tower property of expectation in the previous inequality, one can obtain:
$\mathbb{E}\left[\|x_{T}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{x}_{T}}\|\right]\leq\rho^{T}\|x_{0}-\bar{x}_{0}\|+\frac{\bar{\alpha}\eta\nu\rho^{T}}{\rho+q-1}+\bar{\alpha}C{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\rho^{T-1}}\sum_{k=0}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T-1}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tfrac{1}{(k+1)\log^{2}(k+2)}}.$
Using the Cauchy-Schwarz inequality, and using the fact that $\rho\geq 1-q$ the following can be obtained:
$\mathbb{E}\left[\|x_{T}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{x}_{T}}\|\right]\leq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\rho}^{T}\|x_{0}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{x}_{0}}\|++\frac{\bar{\alpha}\eta\nu\rho^{T}}{\rho+q-1}+{\bar{\alpha}C{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\rho^{T-1}}\sum_{k=0}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T-1}}\left({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tfrac{1}{(k+1)\log^{2}(k+2)}}\right)}.$

### v3

@@ 6 Numerical Experiment
In this experiment, we define $\ell^{\text{val}}_{i}(x_{i})\triangleq\frac{1}{2}\|A_{i}^{\text{val}}x_{i}-b_{i}^{\text{val}}\|^{2},$ where $A_{i}^{\text{val}}\in\mathbb{R}^{n\times d}$ and $b_{i}^{\text{val}}\in\mathbb{R}^{n\times 1}$, and $\ell^{\text{tr}}(x)\triangleq\frac{1}{2}\|A^{\text{tr}}\mathbf{x}-b^{\text{tr}}\|^{2},$ where $A^{\text{tr}}\in\mathbb{R}^{Nn\times Nd}$ and $b^{\text{tr}}\in\mathbb{R}^{Nn\times 1}$ and $X_{i}=\{x_{i}\mid\|x_{i}\|\leq\lambda\}$ for some $\lambda>0$. One can show that this problem can be formulated as (SQVI) by choosing $K(x)=\Pi_{i=1}^{N}K_{i}(x_{(-i)})$, where
$K_{i}(x_{(-i)})=\mbox{argmin}_{x_{i}\in X_{i}}\frac{1}{2}\left\|A_{i}^{\text{tr}}x_{i}+A_{(-i)}^{\text{tr}}x_{(-i)}-b_{i}^{\text{tr}}\right\|^{2},$
@@
Note that, operator $F$ is monotone and satisfies quadratic growth property but it may not be strongly monotone. Furthermore, one can show that Assumption 3 is satisfied (see Theorem 8 in [32] and Corollary 2 in [5]). In Figure 1, we present a performance comparison of our proposed methods. For the triazines dataset [9], we set the number of players $N=6$, for the eunite2001 dataset [9] $N=4$, and for the synthetic dataset $N=10$. In all cases, we utilized $80\%$ of the data points for training and allocated the remaining $20\%$ for validation. To solve the projection inexactly, observe that the sub-problem is a simple bilevel optimization problem. This type of problem has been explored in the literature [21, 44]. Here, following [44], we employed the FISTA algorithm [4] to solve the corresponding regularized problem satisfying Assumption 4. For all the experiments, we execute the inner algorithm for $k\log^{2}(k+1)(1-1\text{e-}3)^{k}$ iterations, and the remaining parameters are selected according to the following table after fine-tuning.
@@ Appendix A Extra Gradient Method
Proof of Corollary 2. (i) By taking expectation from both sides of (5), choosing $N_{k}=N$, and using Assumption 2, the following holds.
@@ Appendix B Gradient Method
Proof of Theorem 2. For any $k\geq 0$, we define $\bar{x}_{k}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\in}\mathcal{P}_{X^{*}}(x_{k})$ where $X^{*}$ denotes the set of optimal solutions of problem (SQVI). From Lemma 2 we conclude that $\bar{x}_{k}=\mathcal{P}_{K(\bar{x}_{k})}[\bar{x}_{k}-\eta F(\bar{x}_{k})]$. Using the update rule of $x_{k+1}$, in Algorithm 2 and the fact that $e_{k}$ denotes the error of computing the projection operator, we obtain the following.
@@
Proof of Corollary 4. Taking expectation from both sides of (10), choosing $N_{k}=\lceil\rho^{-2k}\rceil$, and using Assumption 2, one can obtain:
@@
According to Assumption 4, Algorithm $\mathcal{M}$ has a convergence rate of $C/t_{k}^{2}$ within $t_{k}$ inner steps. By selecting $t_{k}=\tfrac{({k+1})\log^{2}(k+2)}{\rho^{k}}$, we conclude that $\mathbb{E}[\|{e_{k}}\|]\leq\tfrac{C}{t_{k}}=\tfrac{C\rho^{k}}{(k+1)\log^{2}(k+2)}$. By using the tower property of expectation in the previous inequality, one can obtain:
$\mathbb{E}\left[\|x_{T}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{x}_{T}}\|\right]\leq\rho^{T}\|x_{0}-\bar{x}_{0}\|+\frac{\bar{\alpha}\eta\nu\rho^{T}}{\rho+q-1}+\bar{\alpha}C\sum_{k=0}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T-1}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tfrac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\rho^{T-1}}}{(k+1)\log^{2}(k+2)}}.$
@@
Appendix C Almost Sure Convergence
Proof of Theorem 3. By considering $\bar{x}_{k}\in\mathcal{P}_{X^{*}}(x_{k})$ where $X^{*}$ denotes the set of optimal solutions of problem (SQVI) and using Lemma 2 we conclude that $\bar{x}_{k}=\mathcal{P}_{K(\bar{x}_{k})}[\bar{x}_{k}-\eta F(\bar{x}_{k})]$. Now using the update rule of $x_{k+1}$, in Algorithm 2 and denoting $e_{k}$ as the error of computing the projection operator, one can obtain:
$\|x_{k+1}-\bar{x}_{k}\|^{2}$ $=\|(1-\alpha_{k})x_{k}+\alpha_{k}\mathcal{P}_{K(x_{k})}\left[x_{k}-\eta(F(x_{k})+\bar{w}_{k,N_{k}})\right]+{\alpha_{k}e_{k}}-\bar{x}_{k}\|^{2}$
$=\|x_{k}-\bar{x}_{k}-\alpha_{k}(x_{k}-\mathcal{P}_{K(x_{k})}\left[x_{k}-\eta F(x_{k})\right])$
$\quad+\alpha_{k}(e_{k}+\mathcal{P}_{K(x_{k})}\left[x_{k}-\eta(F(x_{k})+\bar{w}_{k,N_{k}})\right]-\mathcal{P}_{K(x_{k})}\left[x_{k}-\eta F(x_{k})\right])\|^{2}.$
Defining $H(x_{k})\triangleq\alpha_{k}(x_{k}-\mathcal{P}_{K(x_{k})}\left[x_{k}-\eta F(x_{k})\right])$, $e^{\prime}_{k}\triangleq e_{k}+\mathcal{P}_{K(x_{k})}\left[x_{k}-\eta(F(x_{k})+\bar{w}_{k,N_{k}})\right]-\mathcal{P}_{K(x_{k})}\left[x_{k}-\eta F(x_{k})\right]$ and using $(a-b+c)^{2}=a^{2}+b^{2}+c^{2}-2ab+2ac-2bc$ we can conclude:
$\|x_{k+1}-\bar{x}_{k}\|^{2}$ $=\|x_{k}-\bar{x}_{k}\|^{2}+\|H(x_{k})\|^{2}+\alpha_{k}^{2}\|e^{\prime}_{k}\|^{2}-2\langle x_{k}-\bar{x}_{k},H(x_{k})\rangle$
$\quad+2\langle x_{k}-\bar{x}_{k},\alpha_{k}e^{\prime}_{k}\rangle-2\langle H(x_{k}),\alpha_{k}e^{\prime}_{k}\rangle.$ (A5)
Furthermore, by using Lemma 1 in the definition of $e^{\prime}_{k}$ and $(a+b)^{2}\leq 2(a^{2}+b^{2})$ one can obtain:
$\|e^{\prime}_{k}\|^{2}$ $\leq 2(\|e_{k}\|^{2}+\|\mathcal{P}_{K(x_{k})}\left[x_{k}-\eta(F(x_{k})+\bar{w}_{k,N_{k}})\right]-\mathcal{P}_{K(x_{k})}\left[x_{k}-\eta F(x_{k})\right]\|^{2})$
$\leq 2(\|e_{k}\|^{2}+\eta^{2}\|\bar{w}_{k,N_{k}}\|^{2}).$ (A6)
On the other hand, since $\bar{x}_{k}\in\mathcal{P}_{X^{*}}(x_{k})$ is a solution of SQVI, one has:
$\langle\eta F(\bar{x}_{k}),\mathcal{P}_{K(x_{k})}\left[x_{k}-\eta F(x_{k})\right]-\bar{x}_{k}\rangle\geq 0.$
Now, let $u=x_{k}-\eta F(x_{k})$ in Lemma 1, since $\bar{x}_{k}\in X$ then the following holds:
$\langle x_{k}-\eta F(x_{k})-\mathcal{P}_{K(x_{k})}\left[x_{k}-\eta F(x_{k})\right],\mathcal{P}_{K(x_{k})}\left[x_{k}-\eta F(x_{k})\right]-\bar{x}_{k}\rangle\geq 0.$
Combining this inequality with the previous one, and using definition of $H(x_{k})$ we get:
$\langle\tfrac{1}{\alpha_{k}}H(x_{k})-\eta(F(x_{k})-F(\bar{x}_{k})),-\tfrac{1}{\alpha_{k}}H(x_{k})+(x_{k}-\bar{x}_{k})\rangle\geq 0.$
Now, rearranging terms and multiplying both sides by $\alpha_{k}$, one can obtain:
$(x_{k}-\bar{x}_{k})^{T}H(x_{k})$ $\geq\tfrac{1}{\alpha_{k}}\|H(x_{k})\|^{2}+\eta\alpha_{k}(x_{k}-\bar{x}_{k})^{T}(F(x_{k})-F(\bar{x}_{k}))-\eta(F(x_{k})-F(\bar{x}_{k}))H(x_{k})$
$\geq\tfrac{1}{\alpha_{k}}\|H(x_{k})\|^{2}+\mu_{F}\eta\alpha_{k}\|x_{k}-\bar{x}_{k}\|^{2}-\tfrac{\eta^{2}L^{2}}{2}\|x_{k}-\bar{x}_{k}\|^{2}-\tfrac{\|H(x_{k})\|^{2}}{2}.$ (A7)
Where the second inequality is obtained by using the QG property, Lipschitz continuity of $F$, and implementing Young’s inequality. Next, substituting (C) and (C) into (C) yields:
$\|x_{k+1}-\bar{x}_{k}\|^{2}$ $\leq\|x_{k}-\bar{x}_{k}\|^{2}+\|H(x_{k})\|^{2}+2\alpha_{k}^{2}(\|e_{k}\|^{2}+\eta^{2}\|\bar{w}_{k,N_{k}}\|^{2})-2\langle x_{k}-\bar{x}_{k},H(x_{k})\rangle$
$\quad+2\langle x_{k}-\bar{x}_{k},\alpha_{k}e^{\prime}_{k}\rangle-2\langle H(x_{k}),\alpha_{k}e^{\prime}_{k}\rangle$
$\leq(1-2\mu_{F}\eta\alpha_{k}+\eta^{2}L^{2}+\alpha_{k}^{2})\|x_{k}-\bar{x}_{k}\|^{2}-(\tfrac{2}{\alpha_{k}}-2)\|H(x_{k})\|^{2}$
$\quad+(4\alpha_{k}^{2}+1)(\|e_{k}\|^{2}+\eta^{2}\|\bar{w}_{k,N_{k}}\|^{2}).$
Now, by considering $\eta=\alpha_{k}$, dropping negative terms and taking the conditional expectation, one can obtain:
$\mathbb{E}[\|x_{k+1}-\bar{x}_{k}\|^{2}]$ $\leq(1+\alpha_{k}^{2}(L^{2}+1))\|x_{k}-\bar{x}_{k}\|^{2}-(\tfrac{2}{\alpha_{k}}-2)\|H(x_{k})\|^{2}$
$\quad+2(2\alpha_{k}^{2}+1)\big{(}(\tfrac{C\rho^{k}}{(k+1)\log^{2}(k+2)})^{2}+\tfrac{\alpha_{k}^{2}\nu^{2}}{N_{k}}\big{)}.$
Where the last term is obtained by knowing the fact that $\mathbb{E}[\|{e_{k}}\|]\leq\tfrac{C}{t_{k}}=\tfrac{C\rho^{k}}{(k+1)\log^{2}(k+2)}$ and Assumption 2. Next, by using the optimality of projection onto a closed convex set and knowing $\bar{x}_{k+1}=\mathcal{P}_{X^{*}}(x_{k+1})$, we can bound the left-hand side of the above inequality, we can conclude:
$\mathbb{E}[\|x_{k+1}-\bar{x}_{k+1}\|^{2}]$ $\leq(1+\alpha_{k}^{2}(L^{2}+1))\|x_{k}-\bar{x}_{k}\|^{2}-(\tfrac{2}{\alpha_{k}}-2)\|H(x_{k})\|^{2}$
$\quad+2(2\alpha_{k}^{2}+1)\big{(}(\tfrac{C\rho^{k}}{(k+1)\log^{2}(k+2)})^{2}+\tfrac{\alpha_{k}^{2}\nu^{2}}{N_{k}}\big{)}.$
Since $\sum_{k=0}^{\infty}\tfrac{1}{\sqrt{N_{k}}}<\infty$, by choosing $\alpha_{k}=\tfrac{1}{k^{0.5+\delta}}$ it is easy to show that the requirements of Lemma 3 are satisfied, and we can conclude that (i) $\|x_{k}-\bar{x}_{k}\|^{2}$ is a convergent sequence. Hence, the sequence $\{x_{k}\}_{k}$ is bounded, which guarantees the existence of a convergent subsequence $x_{k_{q}}\to x^{\#}$ as $q\to\infty$, for some $x^{\#}\in X$. (ii) $\sum_{k=0}^{\infty}(\tfrac{2}{\alpha_{k}}-2)\|H(x_{k})\|^{2}<\infty$ almost surely. From the definition of $H(x_{k})$, it follows that $\sum_{k=0}^{\infty}2\alpha_{k}(x_{k}-\mathcal{P}_{K(x_{k})}[x_{k}-\eta F(x_{k})])<\infty$. Since $\sum_{k=0}^{\infty}\alpha_{k}=\infty$, we conclude that $\lim_{k\to\infty}x_{k}-\mathcal{P}_{K(x_{k})}[x_{k}-\eta F(x_{k})]\to 0$, which in turn implies that $\lim_{q\to\infty}x_{k_{q}}-\mathcal{P}_{K(x_{k_{q}})}[x_{k_{q}}-\eta F(x_{k_{q}})]\to 0$. Now considering (i) and (ii) we can show that $x^{\#}-\mathcal{P}_{K(x^{\#})}[x^{\#}-\eta F(x^{\#})]=0$, which means that $x^{\#}$ is a solution of SQVI. Therefore, we conclude that the sequence $\{x_{k}\}$ has at least one limit point, and every limit point is a solution of the SQVI almost surely. ∎